OpenStudy (anonymous):

Rewrite using the Distributive Property. Then find the product. 3 x 42

OpenStudy (anonymous):

???

OpenStudy (mertsj):

3(40+2)=120+6=126 There you go, smarty

OpenStudy (anonymous):

ooooooooh i was confus ed

OpenStudy (mertsj):

Whatever

OpenStudy (anonymous):

thanks

OpenStudy (mertsj):

yw